pls answer questions (a to c) i will vote you positive . thank you

Assignment 09ACT4105HD - Assignment !! - Deadline: 21 Sep 2020 Email Assignment to tonychuimw@hkbu.edu.hk (in ONE pdf fil
Answer #1

Requirement A

Flexible budget

Out put units

35000

Sales revenue ( 20 * 35000 )

$700000

Less:

Direct materials (12 * 35000 )

$420000

Direct labor ( 2.5 * 35000 )

87500

Indirect labor (5000 + 0.50 * 20000)

22500

Indirect materials

6000

Profit

$164000

Flexible Budget = Budgeted * Actual out put units @ 35000

Budgeted sales price = 800000 / 40000 = 20

Budgeted Direct materials per units = 480000 / 40000 = 12

Direct labor = 100000 / 40000 = 2.5

Indirect labor:- for indirect labor firs we need to calculate variable portion and fixed portion, because it is semi variable cost.

Under high low method

Variable cost per units budgeted = ( Cost @ highest output budgeted - Cost @ lowest output budgeted ) / ( Highest output budgeted - Lowest output budgeted )

Variable cost per units budgeted = ( 25000 - 15000 ) / ( 40000 - 20000 )

Variable cost per units budgeted = 10000 / 20000 = 0.50

Fixed cost total = Total cost at output level - Variable cost per units budgeted * correspondent output level

Fixed cost total = 25000 - 0.50 * 40000 = 25000 - 20000 = 5000

Flexible budget indirect labor cost = 5000 + 0.50 * 35000 = 22500

* Fixed cost are same in budgeted and flexible budget.

Requirement B

Actual

Flexible Budget variance

Flexible

Out put units

35000

35000

Sales revenue

$735000

35000 Favorable

$700000

Less:

Direct materials

$430000

10000 Unfavorable

$420000

Direct labor

80000

7500 Favorable

87500

Indirect labor

31000

8500 Unfavorable

22500

Indirect materials

6500

500 Unfavorable

6000

Profit

$187500

23500 Favorable

$164000

Revenue variance = 35000 Favorable

Total cost variance -= 10000 U + 7500 F + 8500 U + 500 U = 11500 Unfavorable

Profit variance = 23500 favorable

Requirement C

Fixed budget is not flexible and does not change with the actual volume of output achieved. it assumes that criteria would remain static.

It is prepare for one level of sales volume.

Flexible budget is prepare to change according to changed conditions. In flexible budget is prepare for according to several possible output level in relevant range. It actually prepare with actual sales volume.

Importance of flexible budget variance

> The first important is The management can compare actual costs at the actual volume with the budgeted costs at the actual volume.

> these variance can utilize for actual control of cost.

> A flexible budget helps a business to see more variances than a fixed budget.